please show all work, thank you in advance!
Find the exact area of the surface obtained by rotating the curve about the x-axis. 72 y = sin , Os x 5 6 Preview
Answer #1

I Solmis ya sinar, oluto Area Revokie Now surface about aronit Jaryds. where ds = [1 + (not Now, du har = Cosant : ماما Sa jáass site to .(-2) # dt 136 +22 a 4+2 dt s= qxxQ to Sz xox2 1943 dt S=445 7 JEEP + 7 3=47344 +18 sinema] 3=425 135442 + Let 5=